OpenStudy (anonymous):

The function f(x) = 12(4)^x represents the growth of a fly population every year in a remote swamp. Jackie wants to manipulate the formula to an equivalent form that calculates three times a year, not just once a year. Which function is correct for Jackie's purpose, and what is the new growth rate?

OpenStudy (anonymous):

f(x) = 4(4)^x; growth rate 400% f(x) = 4(4)^3x, growth rate 4% f(x) = 4(1.59)^3x; growth rate 159% f(x) = 4(1.59)^x; growth rate 4%